# Deblurring Neural Radiance Fields with Event-driven Bundle Adjustment

### v1

@@ 3.2. Event-driven Bundle Adjustment
In BAD-NeRF (Wang et al., 2023), only the start pose $\mathbf{p}_{1}$ and end pose $\mathbf{p}_{p}$ of the exposure time is parameterized and learned, and the motion process is represented with a linear model. This approach may fail in long exposure or high-speed situations where the camera is likely to have complex motion (non-uniform motion, curved motion) during exposure time. In order to better fit the camera motion trajectory, we express the camera poses $\{\mathbf{P}_{i}\}_{i=1}^{p}$ as $\{\mathbf{T}_{i}\}_{i=1}^{p}\in\mathbf{SE}(3)$.
At this time, in addition to recovering the starting and ending poses, the motion trajectory modeling must also find the relationship between adjacent poses. We can model this correlation by modeling the generation of events, which measures the light intensity change in the log domain asynchronously:
@@
(7) $E(t_{1},t_{2})=\left\{\begin{aligned} \lfloor\frac{\log(I(t_{2}))-\log(I(t_{1})))}{\Theta}\rfloor,&I(t_{2})\leq I(t_{1})\\ \lceil\frac{\log(I(t_{2}))-\log(I(t_{1}))}{\Theta}\rceil,&I(t_{2})>I(t_{1})\end{aligned}.\right.$
Note that we round the modeled events because an event is only triggered when the light intensity change threshold reaches $\Theta$.
According to previous event simulation work, V2E (Hu et al., 2021), we can convert the color of the images into grayscale and model the event generation. Based on this, the estimated event data can be expressed by bringing the predicted color in Eq. (5) into Eq. (7):
(8) $\hat{E}(t_{1},t_{2})=\left\{\begin{aligned} \lfloor\frac{\log(\overline{C}(T({t_{2}})))-\log(\overline{C}(T(t_{1})))}{\Theta}\rfloor,&\overline{C}(t_{2})\leq\overline{C}(t_{1})\\ \lceil\frac{\log(\overline{C}(T({t_{2}})))-\log(\overline{C}(T(t_{1})))}{\Theta}\rceil,&\overline{C}(t_{2})>\overline{C}(t_{1})\end{aligned},\right.$
where $\overline{C}(\cdot)$ means taking the average on the RGB channels to get the intensity value.
The computation of $\hat{E}(\mathbf{x},t_{1},t_{2})$ is differentiable with respect to $\{\mathbf{T}_{i}\}_{i=1}^{p}$, and we can jointly optimize the sampled camera poses and parameters of $F_{\theta}$ by using actual events as supervision.
3.3. Loss Functions
@@
We introduce the intensity-change-metric event loss to optimize the camera motion poses $\{\mathbf{T}_{i}\}_{i=1}^{p}$ during exposure time as well as the NeRF parameter $\theta$ by minimizing the differences between the numbers of predicted and actual events on all pixels $\mathbf{x}\in\mathcal{X}$:
(9) $\mathcal{L}_{event}=\frac{1}{p}\sum_{k=1}^{p-1}\|(\hat{E}_{k}(t_{k},t_{k+1}))-(E_{k}(t_{k},t_{k+1}))\|_{2}^{2}.$
@@
In the training phase, we only calculate the event loss between adjacent sampling poses because if the sampling frames are far apart, the event loss will fluctuate greatly, thus affecting the stability of the network.
@@
We also conduct the photo-metric blur loss between the predicted and the input blurry images:
(10) $\mathcal{L}_{blur}={\|\hat{B}(\mathbf{x})-B(\mathbf{x})\|}_{2}^{2},$
which also has a constraint on both the camera poses and neural radiance fields, as proved in BAD-NeRF. To strengthen the learning, the final loss is the sum of $\mathcal{L}_{event}$ and $\mathcal{L}_{blur}$:
(11) $\mathcal{L}_{blur}=\lambda\mathcal{L}_{event}^{c}+\mathcal{L}_{blur}^{c}+\mathcal{L}_{blur}^{f},$
where $\lambda$ is a weight parameter of event loss. We used the design of the fine and coarse network in NeRF to conduct blur loss $\mathcal{L}^{f}_{blur}$ and $\mathcal{L}^{c}_{blur}$ on both of the networks. Event loss $\mathcal{L}^{f}_{event}$ is only calculated for the fine network because it can render enough texture details for event estimation.
The learnable camera poses $\{\mathbf{T}_{i}\}_{i=1}^{p}$ and NeRF parameter $\theta$ are optimized by the Jacobians of event loss:
(12) $\frac{\partial{\mathcal{L}_{event}}}{\partial\theta}=\frac{\partial\mathcal{L}_{event}}{\partial\mathcal{E}(\mathbf{x})}\cdot\frac{1}{m}\sum_{\mathbf{x\in\mathcal{X}}}\frac{\partial\mathcal{E}(\mathbf{x})}{\partial{C}(\mathbf{x})}\frac{\partial{C}(\mathbf{x})}{\partial\theta},$
(13) $\frac{\partial{\mathcal{L}_{event}}}{\partial\mathbf{T}_{i}}=\frac{\partial\mathcal{L}_{event}}{\partial\mathcal{E}(\mathbf{x})}\cdot\frac{1}{m}\sum_{\mathbf{x\in\mathcal{X}}}\frac{\partial\mathcal{E}(\mathbf{x})}{\partial{C}(\mathbf{x})}\frac{\partial{C}(\mathbf{x})}{\partial\mathbf{T}_{i}},$
and the Jacobians of blur loss:
@@
(14) $\frac{\partial{\mathcal{L}_{blur}}}{\partial\theta}=\frac{\partial\mathcal{L}_{event}}{\partial\mathcal{B}(\mathbf{x})}\cdot\frac{1}{m}\sum_{\mathbf{x\in\mathcal{X}}}\frac{\partial\mathcal{B}(\mathbf{x})}{\partial{C}(\mathbf{x})}\frac{\partial{C}(\mathbf{x})}{\partial\theta},$
(15) $\frac{\partial{\mathcal{L}_{blur}}}{\partial\mathbf{T}_{i}}=\frac{\partial\mathcal{L}_{blur}}{\partial\mathcal{B}(\mathbf{x})}\cdot\frac{1}{m}\sum_{\mathbf{x\in\mathcal{X}}}\frac{\partial\mathcal{B}(\mathbf{x})}{\partial{C}(\mathbf{x})}\frac{\partial{C}(\mathbf{x})}{\partial\mathbf{T}_{i}},$
where $m$ is the number of pixels of the image.
Then, the network can learn an accurate camera motion and sharp 3D implicit representation with gradient propagation.
@@ 3.4. Implementation Details
We train the EBAD-NeRF on a single NVIDIA RTX 3090 GPU. The training time has not increased compared with BAD-NeRF because there is no extra rendering computation for event loss. The sample number of the rays $N=64$ for fine and coarse networks. The number of sampled poses of each view $p$ is set to 4, and the weight of event loss $\lambda$ is set to 0.005. Sec.4.4 evaluates the influence of the parameters. Additionally, we use $\Theta=0.3$ as the threshold of the event generation, which is a typical value in event-based vision (Gallego et al., 2020). A coarse initial pose for each view is given at the start of training.
4. EXPERIEMENTS
@@ 4.1.1. Synthetic Data:
We extend five Blender scenarios of Deblur-NeRF (Cozyroom, Factory, Pool, Tanabata, and Wine). (1) We increase the camera shake amplitude of the Factory, Pool, Tanabata, and Wine scenes to increase the blur degree of the training images and maintain the camera shaking in Deblur-NeRF of the Cozyroom scene to evaluate the effect of different shaking scales. (2) During the camera motion, we added the speed changes often occurring when the hand-held camera shakes. (3) To generate corresponding event data, we input the virtual sharp frames of the camera shake process rendered by Blender into the event simulation tool V2E (Hu et al., 2021) to generate the corresponding event stream.
@@ 4.1.2. Real Data:
To test the effectiveness of our method on real captured data, we captured two sets of real data using the DAVIS-346 (Taverni et al., 2018) event camera, which can capture spatial-temporal aligned event data and RGB data at the same time. Both sets of data are captured in low-light scenes, so a longer exposure time (100ms) is needed for RGB data to get a bright enough image. We use a handheld camera to capture training data with blur and a tripod to mount the camera to capture sharp testing data. Each scene contains 16 views of blurry images and corresponding events for training and 4 novel view sharp images for training.
4.2. Comparing Method
4.2.1. Image Deblurring Method:
For comparison, we selected two classic learning-based image deblurring methods, MPR (Zamir et al., 2021a) and SRN (Tao et al., 2018b). In order to make a more fair comparison in terms of data modalities, we selected the D2Net method, which also uses event and image modal data fusion to achieve image deblurring. We input the images deblurred by these three methods into the original NeRF for the novel view generation task to learn a sharp 3D representation.
4.2.2. Deblurring NeRF Method:
We selected image-based deblurring NeRF method Deblur-NeRF (Ma et al., 2022) and BAD-NeRF (Wang et al., 2023) and ERGB-dual-modal-based deblurring NeRF method E2NeRF (Qi et al., 2023) for comparison. For Deblur-NeRF, we use the default parameters for the spare deformation kernel. For BAD-NeRF, we use the number of interpolated virtual cameras as 7 to get sufficiently satisfactory results, as described in the paper. For E2NeRF, we follow the paper and use $b=4$ to train the model. All NeRF-based methods are trained with 5000 rays of batch size and 64 sampling points for coarse and fine networks. We perform 100,000 iterations on synthetic data. On real data, due to the reduction in image resolution and the number of views, 50,000 iterations are enough for all NeRF-based methods to converge to good enough results.
@@ 4.2.3. Metrics:
We use PSNR, SSIM, and LPIPS (Zhang et al., 2018) to evaluate the NeRF reconstruction results and use absolute trajectory error (ATE) to evaluate the fitting accuracy of the camera motion trajectory (Grupp, 2017).
@@
PSNR↑ SSIM↑ LPIPS↓ ATE
@@
Figure 4. Ablation study of trajectory fitting accuracy on Tanabata scene. EBAD-NeRF-linear and EBAD-NeRF-noe are defined in Sec. 4.3. The figure shows that the event-driven bundle adjustment significantly fits the camera motion closer to the ground truth in the roll, pitch, and yaw metrics than other methods.
@@
EBAD-NeRF 30.53 .9475 .1120 28.10 .9085 0.159 31.50 .9193 .1607 24.91 .8783 .2071 26.66 .8732 .2230 28.30 .9031 .1749
4.3. Event-driven Bundle Adjustment Analysis
We analyze the proposed event-driven bundle adjustment at two levels: 3D reconstruction quality and motion trajectory fitting accuracy. As shown in Table 2, Figure 3 and Figure 4, EBAD-NeRF-noe training proposed EBAD-NeRF without event data as supervision. EBAD-NeRF-linear represents the results of adding event data constraints to BAD-NeRF directly.
4.3.1. Defect of Linear Interpolation Bundle Adjustment:
The quantitative results in Table 2 demonstrate that without event enhancement, the bundle adjustment constrained by linear interpolation (BAD-NeRF) is even worse than the unconstrained bundle adjustment (EBAD-NeRF-noe) for fitting irregular camera motions. The same conclusion will be obtained by comparing EBAD-NeRF-linear and EBAD-NeRF-full, both with event enhancement.
4.3.2. Effect of Linear Event in Bundle Adjustment:
Comparing BAD-NeRF and EBAD-NeRF-linear, even with the negative impact of linear interpolation bundle adjustment in Table 2, EBAD-NeRF-linear still has improvements in trajectory fitting and reconstruction results. Comparing EBAD-NeRF-linear and EBAD-NeRF under the constraints of event information, event-driven bundle adjustment achieved the best trajectory fitting and reconstruction results.
To sum up, during the NeRF training, event data can not only provide image-level optimization information but also pose-level practical constraints. Our proposed EBAD-NeRF effectively utilizes the ERGB-dual-modal data to achieve this goal. The qualitative comparison in Figure 3 and Figure 4 also testify to this.
Figure 5. The influence of the number of learnable poses $p$ and the weight parameter $\lambda$ on the Cozyroom scene. The results are average for reconstructed images from both deblurring and novel views.
4.4. Analysis on Parameters
We evaluate the rendering results on the Cozyroom scene to analyze the superparameters $p$ and $\lambda$ of the proposed network. The red, green, and blue lines in Figure 5 represent PSNR, SSIM, and LPIPS, respectively.
4.4.1. The influence of $b$:
Intuitively, the more virtual sharp frames, the more accurate the simulation of camera motion blur, and the corresponding network learning results will be better. The results in Figure 5 are consistent with this. When the number of virtual frames gradually increases from 1, the performance of the network also becomes significantly better. After exceeding 5, the number of virtual frames has almost no impact on the reconstruction results. Therefore, $p=5$ is set in our experiments.
4.4.2. The influence of $\lambda$:
As shown in Figure 5, $\lambda$ has the best effect when set to 0.005, and if the weight of event loss is further increased, the network performance will gradually decrease.
@@ 4.4.2. The influence of $\lambda$:
NeRF NeRF+D2Net NeRF+MPR NeRF+SRN Deblur-NeRF BAD-NeRF E2NeRF EBAD-NeRF
PSNR↑ 24.87 26.31 24.35 25.91 25.59 28.10 27.78 29.60
SSIM↑ .8562 .8829 .8539 .8706 .8744 .8839 .9062 .9175
LPIPS↓ .5005 .4092 .4963 .4198 .3885 .3632 .2403 .2156

### v2

@@ 3.2. Event-driven Bundle Adjustment
In long exposure or high-speed situations, the camera is likely to have complex motion (non-uniform motion) during exposure time. To better model the motion blur, we transform the sampled $p$ poses $\{\mathbf{P}_{i}\}_{i=1}^{p}$ into $p$ learnable pose $\{\mathbf{T}_{i}\}_{i=1}^{p}\in\mathbf{SE}(3)$ at time $\{t_{i}\}_{i=1}^{p}$. Then, we can use events as a bridge between camera pose changes and neural radiance fields to bundle these poses into actual camera motion trajectory during the NeRF optimization.
As described in Sec.3.1, we can express the virtual sharp image color during the exposure as $C(\mathbf{T}_{i},\mathbf{x})=C(\mathbf{P}_{i},\mathbf{x})$ and convert it into grayscale as $G(\mathbf{T}_{i},\mathbf{x})$ by averaging the RGB channels. Then, according to previous event simulation work (Hu et al., 2021), we transform $G(\mathbf{T}_{i},\mathbf{x})$ into the log function domain as Ł$\L(\mathbf{T}_{i},\mathbf{x})$ and simulate the generation of events caused by camera motion as:
@@
(7) ^$\hat{\mathcal{E}}(\mathbf{x},t_{i},t_{i+1})=\left\{\begin{aligned} \frac{\L(\mathbf{T}_{i+1},\mathbf{x})-\L(\mathbf{T}_{i},\mathbf{x})}{\Theta},&\L(\mathbf{T}_{i+1},\mathbf{x})\leq\L(\mathbf{T}_{i},\mathbf{x})\\ \frac{\L(\mathbf{T}_{i+1},\mathbf{x})-\L(\mathbf{T}_{i},\mathbf{x})}{\Theta},&\L(\mathbf{T}_{i+1},\mathbf{x})>\L(\mathbf{T}_{i},\mathbf{x})\end{aligned}.\right.$
@@
Then with the real events $\mathcal{E}(\mathbf{x},t_{i},t_{i+1})$ captured by the event camera as event-driven bundle adjustment, we can optimize adjunctive poses $T_{i}$ and $T_{i+1}$ with the intensity-change-metric event loss by minimizing the differences between the numbers of predicted and actual events on all $m$ pixels $\mathbf{x}\in\mathcal{X}$ among all input views:
(8) $\mathcal{L}_{event}=\frac{1}{m}\sum_{\mathbf{x}\in\mathcal{X}}\frac{1}{p-1}\sum_{i=1}^{p-1}\|(\hat{\mathcal{E}}(\mathbf{x},t_{i},t_{i+1}))-(\mathcal{E}(\mathbf{x},t_{i},t_{i+1}))\|_{2}^{2}.$
The computation of ^$\hat{\mathcal{E}}(\mathbf{x},t_{1},t_{2})$ is differentiable with respect to $\{\mathbf{T}_{i}\}_{i=1}^{p}$ and latent sharp NeRF parameters $\theta$, which is optimized by the Jacobians of event loss:
(9) $\frac{\partial{\mathcal{L}_{event}}}{\partial\theta}=\frac{\partial\mathcal{L}_{event}}{\partial\hat{\mathcal{E}}(\mathbf{x})}\cdot\frac{1}{m}\sum_{\mathbf{x\in\mathcal{X}}}\frac{\partial\hat{\mathcal{E}}(\mathbf{x})}{\partial{C}(\mathbf{x})}\frac{\partial{C}(\mathbf{x})}{\partial\theta},$
(10) $\frac{\partial{\mathcal{L}_{event}}}{\partial\mathbf{T}_{i}}=\frac{\partial\mathcal{L}_{event}}{\partial\hat{\mathcal{E}}(\mathbf{x})}\cdot\frac{1}{m}\sum_{\mathbf{x\in\mathcal{X}}}\frac{\partial\hat{\mathcal{E}}(\mathbf{x})}{\partial{C}(\mathbf{x})}\frac{\partial{C}(\mathbf{x})}{\partial\mathbf{T}_{i}},$
@@
Notice that in Eq. (8) we only calculate the event loss between adjacent poses because if the sampled poses are too far apart, the event loss will fluctuate greatly, reducing the training stability.
3.3. Final Loss
@@
We conduct the photo-metric blur loss between the predicted and the input blurry images as in NeRF for all $m$ pixels $\mathbf{x}\in\mathcal{X}$ among all input views:
(11) $\mathcal{L}_{blur}=\frac{1}{m}\sum_{\mathbf{x}\in\mathcal{X}}{\|\hat{\mathcal{B}}(\mathbf{x})-\mathcal{B}(\mathbf{x})\|}_{2}^{2},$
The blur loss not only has a constraint on camera poses $\{\mathbf{T}_{i}\}_{i=1}^{p}$ to avoid the spreading of poses but also plays a main role in supervising NeRF parameters $\theta$ to learn texture details of the scene that event data cannot provide:
@@
(12) $\frac{\partial{\mathcal{L}_{blur}}}{\partial\theta}=\frac{\partial\mathcal{L}_{event}}{\partial\hat{\mathcal{B}}(\mathbf{x})}\cdot\frac{1}{m}\sum_{\mathbf{x\in\mathcal{X}}}\frac{\partial\hat{\mathcal{B}}(\mathbf{x})}{\partial{C}(\mathbf{x})}\frac{\partial{C}(\mathbf{x})}{\partial\theta},$
(13) $\frac{\partial{\mathcal{L}_{blur}}}{\partial\mathbf{T}_{i}}=\frac{\partial\mathcal{L}_{blur}}{\partial\hat{\mathcal{B}}(\mathbf{x})}\cdot\frac{1}{m}\sum_{\mathbf{x\in\mathcal{X}}}\frac{\partial\hat{\mathcal{B}}(\mathbf{x})}{\partial{C}(\mathbf{x})}\frac{\partial{C}(\mathbf{x})}{\partial\mathbf{T}_{i}}.$
The final loss is defined as:
(14) $\mathcal{L}_{blur}=\lambda\mathcal{L}_{event}^{f}+\mathcal{L}_{blur}^{c}+\mathcal{L}_{blur}^{f},$
where $\lambda$ is a weight parameter of event loss. We used the design of the fine and coarse network in NeRF and conduct blur loss $\mathcal{L}^{f}_{blur}$ and $\mathcal{L}^{c}_{blur}$ on both of the networks. Event loss $\mathcal{L}^{f}_{event}$ is only calculated for the fine network because it can render enough texture details for precise event estimation. Then, the network can learn an accurate camera motion and latent sharp 3D implicit representation with gradient propagation.
@@ 3.4. Implementation Details
We train the EBAD-NeRF on a single NVIDIA RTX 3090 GPU. The training time is close to BAD-NeRF. The sample number of the rays is $N=64$ for both fine and coarse networks. The number of sampled poses of each view $p$ is set to 5, and the weight of event loss $\lambda$ is set to 0.005. Sec.4.4 evaluates the influence of the parameters. Additionally, we use $\Theta=0.3$ as the threshold of the event generation, which is a typical value in event-based vision (Gallego et al., 2020). A coarse initial pose for each view is given at the start of training.
4. Experiments
@@ 4.1.1. Synthetic Data:
We use the five Blender scenarios of Deblur-NeRF (Cozyroom, Factory, Pool, Tanabata, and Wine) to generate our synthetic data. We increase the camera shake amplitude to generate a severe blur and add the camera motion speed change during the exposure to simulate a non-uniform camera motion blur for the Factory, Pool, Tanabata, and Wine scenes. Besides, we input the virtual sharp frames during the camera shake process rendered by Blender into the V2E (Hu et al., 2021) to generate the corresponding event stream as in the synthetic datasets of E2NeRF.
@@ 4.1.2. Real Data:
To test the effectiveness of our method on real data, we capture two sets of real data (Bar and Classroom) with DAVIS-346 (Taverni et al., 2018) event camera, which can capture spatial-temporal aligned event data and RGB data at the same time. Both sets of data are captured in low-light scenes, so RGB data requires a longer exposure time (100ms) to get a bright enough image. We use a handheld camera to capture blurry images and events for training and a tripod to capture sharp images for testing. Each scene contains 16 views of blurry images and corresponding events for training and 4 novel view sharp images for testing.
4.2. Comparing Methods and Metrics
4.2.1. Image Deblurring Methods
For comparison, we selected two classic learning-based image deblurring methods, MPR (Zamir et al., 2021a) and SRN (Tao et al., 2018b). In order to make a more fair comparison in terms of data modalities, we use D2Net (Shang et al., 2021), which also uses events and blurry image data to achieve image deblurring. We input the images deblurred by these three methods into the original NeRF for the novel view generation task and named them as MPR+NeRF, SRN+NeRF, and D2Net+NeRF.
4.2.2. Deblurring NeRF Methods:
We selected image-based deblurring NeRF methods Deblur-NeRF (Ma et al., 2022) and BAD-NeRF (Wang et al., 2023) and ERGB deblurring NeRF method E2NeRF (Qi et al., 2023). All NeRF-based methods are trained with 5000 rays of batch size and 64 sampling points for coarse and fine networks. Other parameters are the default values of the methods. We perform 100,000 iterations on synthetic data. On real data, due to the reduction in image resolution and the number of input views, 50,000 iterations are enough for all NeRF-based methods to converge to the final results.
@@ 4.2.3. Metrics:
We use PSNR, SSIM, and LPIPS (Zhang et al., 2018) to evaluate the reconstruction results and use absolute trajectory error (ATE) to evaluate the fitting accuracy of the camera motion trajectory (Grupp, 2017).
@@
PSNR↑ SSIM↑ LPIPS↓ ATE↓
@@
Figure 4. Qualitative ablation study of trajectory fitting accuracy on Tanabata scene. EBAD-NeRF-linear and EBAD-NeRF-noe are defined in Sec. 4.3. EBAD-NeRF significantly fits the camera motion closer to the ground truth in the roll, pitch, and yaw metrics than all other methods.
4.3. Event-driven Bundle Adjustment Analysis
We analyze the proposed event-driven bundle adjustment at two levels: 3D reconstruction quality and motion trajectory fitting accuracy. In Table 2 and Figures 3, 4, EBAD-NeRF-noe training the proposed EBAD-NeRF without event data as supervision and EBAD-NeRF-linear represents the results of adding event data constraints to BAD-NeRF with linear interpolation directly. Results in Table 2 are the average of the deblurring view and novel view.
4.3.1. Defect of Linear Interpolation Bundle Adjustment:
The quantitative results in Table 2 demonstrate that without event enhancement, the bundle adjustment constrained by linear interpolation (BAD-NeRF) is even worse than the unconstrained bundle adjustment (EBAD-NeRF-noe) when facing complex camera motion. With event enhancement, the same conclusion can be obtained by comparing EBAD-NeRF-linear and EBAD-NeRF-full.
4.3.2. Effect of Event in Bundle Adjustment:
Comparing BAD-NeRF and EBAD-NeRF-linear, even with the negative impact of linear pose interpolation, EBAD-NeRF-linear still has improvements in trajectory fitting and reconstruction results with event enhancement as shown in Table 2. Comparing EBAD-NeRF-noe and EBAD-NeRF, introducing events significantly improves the accuracy of trajectory estimation and reconstruction results.
To sum up, during the NeRF training, event data can not only provide image-level optimization information but also pose-level practical constraints. Our proposed EBAD-NeRF effectively utilizes the ERGB-dual-modal data to achieve this goal. The qualitative comparison in Figure 3 and Figure 4 also testify to this, where EBAD-NeRF generates the sharpest rendering results and closest trajectory estimation to the ground truth.
4.4. Parameters Analysis
We evaluate the rendering results of the Cozyroom scene to analyze the influence of $p$ and $\lambda$ of our method in Figure 5.
4.4.1. The influence of $p$:
Intuitively, the more virtual sharp frames, the more accurate the simulation of camera motion blur will be, and the corresponding network learning results will be better. The results in Figure 5 are consistent with this. When the number of virtual frames gradually increases from 1, the network’s performance also significantly improves. However, after exceeding 5, the number of virtual frames has almost no impact on the reconstruction results, and more virtual sharp frames will require more calculations. Therefore, $p=5$ is set in our experiments.
4.4.2. The influence of $\lambda$:
As shown in Figure 5, when we take $\lambda$ as 0.005, EBAD-NeRF achieves the best results on the three metrics.
Figure 5. Evaluation on the influence of the number of the sampled poses $p$ and the weight parameter $\lambda$. The results are averages of reconstructed images on both deblurring and novel views of the Cozyroom scene. The red, green, and blue lines represent PSNR, SSIM, and LPIPS, respectively.
Table 3. Absolute Trajectory Error Results of Blender Scenes.
ATE↓
COLMAP-Blur (NeRF (Mildenhall et al., 2021)) 0.0476 $\pm$ 0.0171
COLMAP-EDI (E2NeRF (Qi et al., 2023)) 0.0419 $\pm$ 0.0159
BAD-NeRF (Linear Interpolation(Wang et al., 2023)) 0.0462 $\pm$ 0.0142
BAD-NeRF (Cubic Interpolation(Wang et al., 2023)) 0.0542 $\pm$ 0.0207
Event-driven Bundle Adjustment (Ours) 0.0383 $\pm$ 0.0143
@@ 4.4.2. The influence of $\lambda$:
EBAD-NeRF 30.53 .9475 .1120 28.10 .9085 .1717 31.50 .9193 .1607 24.91 .8783 .2071 26.66 .8732 .2230 28.34 .9054 .1749
@@
NeRF NeRF+D2Net NeRF+MPR NeRF+SRN Deblur-NeRF BAD-NeRF BAD-NeRF-Cubic E2NeRF EBAD-NeRF
PSNR↑ 24.87 26.31 24.35 25.91 25.59 28.10 27.03 27.78 29.60
SSIM↑ .8562 .8829 .8539 .8706 .8744 .8839 .8744 .9062 .9175
LPIPS↓ .5005 .4092 .4963 .4198 .3885 .3632 .2978 .2403 .2156